if the sum of two consecutive odd integers is divided by 4, the quotient is 10. find the numbers.
Answer by mananth(12270)

Let the consequtive odd numbers be x & x-2
x+x-2 / 4 = 10
2x-2 = 40
2x=40+2
2x=42
x=21
the other number = 19.
Check
19+21 / 4 = 10